OpenStudy (gabylovesyou):

Part 1: Solve the inequality -5x + 8

OpenStudy (pradipgr817):

So we need to solve it together

OpenStudy (gabylovesyou):

alright

OpenStudy (gabylovesyou):

x >/= 2

OpenStudy (gabylovesyou):

but how do i put it step by step??

OpenStudy (pradipgr817):

takes Xs and constant on either sides

OpenStudy (pradipgr817):

8+4-5x <=x 12<=6x 2<=x

OpenStudy (pradipgr817):

so it will be a closed circle on 2 and shading on left

OpenStudy (gabylovesyou):

@pradipgr817 write it like this: Part 1: Part 2: so i understand :p

OpenStudy (pradipgr817):

part 1 takes Xs and constant on either sides 8+4-5x <=x 12<=6x 2<=x

OpenStudy (pradipgr817):

Part 2: so it will be a closed circle on 2 and shading on left

OpenStudy (gabylovesyou):

Does 2<=x equal the same as x >/= 2

OpenStudy (pradipgr817):

you can add a simplification step 8+4-5x <=x 8+4<=x+5x 12<=6x 12/6<=x 2<=x

OpenStudy (gabylovesyou):

for part 1?

OpenStudy (pradipgr817):

yea, the legs of the symbol opening toward X

OpenStudy (gabylovesyou):

alrighty then...............

OpenStudy (gabylovesyou):

i added it to part 1

OpenStudy (pradipgr817):

Is it clear?

OpenStudy (anonymous):

we have -5x+8<=x-4 so we add 5x to each side of our inequality andwipe out that awkward -5x...we now have 8<=6x-4...now wipe out the -4 on right hand side by adding 4 to each side.. we are left with 12,<=6x..so we divide each side by 6 and get 2<=x or we can say x=>2 This is represented on a no. line by a shaded in circle above the number 2 and a line drawn to the right of this circle

OpenStudy (gabylovesyou):

yes thanks

OpenStudy (gabylovesyou):

@solutions Thank you sooo much :)